solve plz? with rules of inference and replacement...
1. (~D + A) v (~D + R)
2. D v ~A /∴ K > R
_______________________________
1. ~(G + M)
2. M v ~G /∴ ~G
_______________________________
I know y'all only answer two questions...but it's worth a shot to ask...If not, thank you anyways!!
1. P = (R + S)
2. P /∴ R = S

Answer by math_tutor2020(3816)


Normally I would mention that the rule of this website is "one question per post", but the problems aren't too lengthy here.

Problem 1
NumberStatementLine(s) UsedReason
1(~D + A) v (~D + R)
2D v ~A
:.K > R
3~D + (A v R)1Distribution
4~D3Simplification
5~A2,4Disjunctive Syllogism
6A v R3Simplification
7R6,5Disjunctive Syllogism
8R v ~K7Addition
9~K v R8Commutation
10K > R9Material Implication

Refer to these rules of inference and replacement
https://logiccurriculum.com/2019/02/09/rules-for-proofs/
The notation is slightly different. That reference sheet has a center dot instead of a plus sign. Also, that sheet uses a horsehoe instead of a greater than sign.

==================================================================================================================
Problem 2

NumberStatementLine(s) UsedReason
1~(G + M)
2M v ~G
:.~G
3~G v ~M1De Morgan's Law
4~G v M2Commutation
5G > ~M3Material Implication
6G > M4Material Implication
7~M > ~G6Transposition
8G > ~G5,7Hypothetical Syllogism
9~G v ~G8Material Implication
10~G9Tautology


==================================================================================================================
Problem 3

This logical argument is invalid

Proof of this is shown in the truth table below.
PremisePremiseConclusion
PRSR+SP = (R+S)PR = S
TTTTTTT
TTFTTTF
TFTTTTF
TFFFFTT
FTTTFFT
FTFTFFF
FFTTFFF
FFFFTFT

Here is a review of various truth table rules
https://www.algebra.com/algebra/homework/Conjunction/truth-table1.lesson

Focus on row 2. I've highlighted this in red
The premises P = (R+S) and P are true, but the conclusion is false.
Any time there is a situation with all true premises but a false conclusion, it means the argument is invalid.

Therefore, you would not be able to find a logical derivation when presented with these premises and conclusion. Attempting to do so will have you endlessly searching for something that doesn't exist. It's possible that your teacher made a typo somewhere.